Find the basis function of the differential equation using Frobenius method

2ax(1 y (1-5x)-y = 0
Answer #1

2x (1-x)y+ 1-5x)y-y= o Let y=Cm be a trial olの Q = y = MCm y m/m-1) CmY putly y,リ, y Jn gm lwe get y =0 11-5x)y! 2x(1-x)y2m+)m Cm 二0 -15mcn- -1am 2mm-+5m m m jatry O +3M+1 Camt Cm Basir bunction Ns our Soln i y=Cmm 0-u y=Co+Cx+ C2+Cg+y x+ .... Cm